# Learning Enhanced Ensemble Filters

### v1

@@ 5.4 Lorenz ’63
Obs Time Step $\Delta t=0.15$
(a) $\sigma_{y}=0.7$
(b) $\sigma_{y}=1$
(c) Relative improvement
Figure 3: Comparison results on the Lorenz ’63 system. Panels (a) and (b) show direct RMSE comparisons between different methods at observation noise levels $\sigma_{y}=0.7$ and $\sigma_{y}=1$, respectively. Panel (c) illustrates the relative improvement of our fine-tuning method compared to benchmarks. Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain).
@@
In Figure 3, we compare the performance on R-RMSE (5.3) of our pretrained model and fine-tuned model against four benchmark methods: EnKF, ESRF, and IEnKF, for the Lorenz ’63 system with $\sigma_{y}=0.7$ (Figure 3(a)) and $\sigma_{y}=1.0$ (Figure 3(b)). The LETKF method is not included in this comparison since the system has only three dimensions, making localization unnecessary. For a fair comparison, we turn off the localization part in our architecture by setting all localization weights to be $1$. We only fine-tune the parameters $\theta_{\mathrm{gain}}$ for $F^{\mathrm{gain}}$ (4.28) and $\theta_{\mathrm{infl}}$ for $F^{\mathrm{infl}}$ (4.31) for ensemble sizes $N\neq 10$. In Figure 3(c), we show the relative improvement $\mathcal{E}_{\mathrm{RI}}$ (5.4) of our fine-tuned method compared to EnKF and IEnKF.
@@
The IEnKF [15] uses the same observational information as the standard EnKF, but reformulates the filtering problem as a lag-1 smoothing problem, and solves it iteratively to better handle nonlinearities. While the IEnKF demonstrates excellent performance in low-dimensional systems (e.g., the Lorenz ’63 model, $d=3$), its effectiveness diminishes as dimensionality increases (e.g., the Lorenz ’96 model, $d=40$).
This deterioration occurs due to lack of localization in the implementation of the IEnKF we use; however, even with large ensembles with less need for localization, our method outperforms the IEnKF in the higher-dimensional systems. Moreover, IEnKF generally provides a substantial improvement over non-iterative methods only when the observations are sufficiently dense, as shown in [15] using experiments with the Lorenz ’96 model.
@@ 5.5 Robustness To Inherent Randomness
Due to the inherent randomness, the same method may exhibit varying performance across different test trajectories. To mitigate this randomness and obtain a more reliable assessment of the method’s effectiveness, we evaluate the performance on $M_{\mathrm{test}}$ trajectories and report the mean R-RMSE $\bar{\mathcal{E}}$ (5.3). While the mean R-RMSE provides a measure of overall accuracy, we are interested in the stability of the method across different trajectories. To quantify this stability, we compute the standard deviation (STD) of the R-RMSE values across all test trajectories $\{V_{m}^{\dagger}\}_{m=1}^{M_{\mathrm{test}}}$ and their corresponding estimated trajectories $\{V_{m}\}_{m=1}^{M_{\mathrm{test}}}$, defined as:
$\sigma_{\mathcal{E}}=\sqrt{\frac{1}{M_{\mathrm{test}}}\sum_{m=1}^{M_{\mathrm{test}}}\left(\mathcal{E}(V_{m},V_{m}^{\dagger})-\bar{\mathcal{E}}\right)^{2}}$ (5.8)
@@
Table 4: Standard deviation (STD) of the relative root mean square error (R-RMSE) as defined in (5.8). We consider both Pretrain* and Tuned* variants of our MNMEF method.
A smaller STD indicates better stability of the method. The lowest STD in each column is highlighted in bold. Method Lorenz ’96 (5.5) KS (5.6) Lorenz ’63 (5.7) $\sigma_{y}=1.0$ $\sigma_{y}=0.7$ $\sigma_{y}=1.0$ $\sigma_{y}=0.7$ $\sigma_{y}=1.0$ $\sigma_{y}=0.7$ Pretrain* 1.03e-2 1.26e-2 8.60e-3 1.07e-2 1.78e-3 1.40e-3 Tuned* 1.08e-2 1.12e-2 8.91e-3 1.07e-2 1.79e-3 1.29e-3 EnKF [8] 2.16e-2 3.10e-2 2.66e-2 3.15e-2 3.41e-3 2.32e-3 ESRF [47] 3.09e-2 1.95e-2 2.59e-2 3.27e-2 2.78e-3 2.01e-3 LETKF [14] 2.63e-2 6.13e-2 2.66e-2 3.58e-2 - - IEnKF [15] 1.83e-2 2.61e-2 2.23e-2 3.03e-2 2.15e-3 1.44e-3
@@
It is worth noting that in data assimilation literature, there exist metrics similar to variance or standard deviation that are used to measure the spread of ensemble-based methods [6]. However, since our focus in this paper is on state estimation performance, we are not concerned with spread metrics. The stability measure $\sigma_{\mathcal{E}}$ defined in (5.8) is fundamentally different from spread metrics as it quantifies the consistency of estimation errors across different test trajectories rather than the dispersion of ensemble members.
Based on Table 4, we observe that our pretrained MNMEF and fine-tuned MNMEF demonstrate comparable standard deviation values, both of which are significantly lower than all benchmark methods. This indicates that our approaches maintain more consistent performance across different test trajectories. The enhanced stability suggests that our methods are more reliable in practical applications.
@@ 5.6 Inflation and Localization
In this section, we will conduct further experiments on the learning inflation (Subsection 4.2.2) and localization (Subsection 4.2.3) components from the architecture of our MNMEF method. These experiments demonstrate that what our method learns has many similarities with what is used in the application of inflation and localization techniques to the benchmark methods.
@@ 5.6.1 Inflation Experiments
In Subsection 4.2.2, we introduce our learning-based inflation scheme (4.30) which depends on a learned term $\widehat{u}^{(n)}_{\theta}$ output from the neural network $F^{\mathrm{infl}}$ (4.31).
(a) $\sigma_{y}=0.7$
(b) $\sigma_{y}=1$
(c) Relative Difference
Figure 4: Comparison of the trained inflation versus the no inflation scenario when applied to the KS system (5.6), using our MNMEF method, fine-tuned for different ensemble sizes. Panels (a) and (b) compare RMSE with and without inflation at observation noise levels $\sigma_{y}=0.7$ and $\sigma_{y}=1$, respectively. Panel (c) shows the relative performance difference (5.9) between these scenarios.
@@
For our experimental approach, we examine models that are first pretrained and then fine-tuned separately for various ensemble sizes. During the inference process of these models, we manually set the inflation term to zero, i.e. $\widehat{u}^{(n)}_{\theta}=0$, and analyze the resulting changes in the average R-RMSE $\bar{\mathcal{E}}$ (5.3). We conduct the experiments on the KS dynamical system (5.6) only.
Our results are presented in Figure 4. We compare the RMSE $\mathcal{E}$ between our standard architecture with the trained inflation (blue curve) and the no inflation version by setting $\widehat{u}^{(n)}_{\theta}=0$ (red curve) for observation noise $\sigma_{y}=0.7$ (Figure 4(a)) and $\sigma_{y}=1$ (Figure 4(b)). In addition, we illustrate the relative difference in Figure 4(c) define as:
$\mathcal{E}_{r}=\frac{\mathcal{E}(\mathrm{NoInfl})-\mathcal{E}(\mathrm{TrainedInfl})}{\mathcal{E}(\mathrm{TrainedInfl})}$ (5.9)
The results in Figure 4 clearly demonstrate the beneficial impact of our trained inflation approach. Figure 4(c) illustrates the relative difference metric $\mathcal{E}_{r}$, which consistently shows positive values across all assimilation steps, indicating that removing the inflation term ($\widehat{u}^{(n)}_{\theta}=0$) degrades performance.
In addition, our method remains functional with increased RMSE values when we remove inflation by setting $\widehat{u}^{(n)}_{\theta}=0$ (equivalent to $\alpha=1$ in EnKF). This addresses our concern, because, if inflation were dominating the architecture by masking the contribution of the state $v^{(n)}$, removing it would cause model collapse.
@@ 5.6.2 Localization Experiments
In this subsection, we present experimental comparisons between our learned function $g_{\theta}$ and the Gaspari–Cohn (GC) function with the optimal localization radius, found by grid-search, for LETKF – see Figure 5 . The optimal GC radius values chosen are 1, 3, and 5 (multiplied by $\sqrt{10/3}$ when implemented) for ensemble sizes of 10, 20, and 100, respectively. The learned function $g_{\theta}$ (shown as red curves, mean $\pm$ 1 std across time steps due to adaptivity) demonstrates clear differences compared to the fixed GC function (blue curves). Notably, our approach dynamically adjusts the localization weights at each assimilation step based on ensemble distribution, while the GC function remains static throughout the assimilation process.
@@
The results in Figure 5 demostrates the effectiveness of our learned localization. Despite imposing no explicit constraints on shape or monotonicity, the learned function $g_{\theta}$ naturally exhibits similar behavior to the GC function: it decreases monotonically with distance and shares a similar shape. A key difference is that our function occasionally exceeds values of 1 at short distances, reflecting a combined localization and inflation effect where covariance between closely spaced dimensions is amplified, since we are learning the inflation and localization simutaniously. This adaptivity allows $g_{\theta}$ to dynamically adjust to the ensemble distribution, leading to improved performance compared to the fixed GC function, as supported by RMSE results in Subsections 5.45.3.
@@ 5.7 Fine-Tuning for Different Ensemble Size
Firstly, we consider the time complexity between the pretraining stage and fine-tuning stage. Let $P_{t}$ and $P_{ft}$ denote the number of total parameters and fine-tuning parameters respectively, i.e.
$P_{t}$ $=|\theta_{\mathrm{ST}}|+|\theta_{\mathrm{gain}}|+|\theta_{\mathrm{infl}}|+|\theta_{\mathrm{loc}}|,$ (5.10a)
$P_{ft}$ $=|\theta_{\mathrm{gain}}|+|\theta_{\mathrm{infl}}|+|\theta_{\mathrm{loc}}|.$ (5.10b)
For a fixed dynamic and training trajectories, the time complexity for pretraining with the ensemble size $N$ is $\mathcal{O}\bigl{(}ENP_{t}\bigr{)}$ while the fine-tuning with the ensemble size $N^{\prime}$ has the time complexity $\mathcal{O}\bigl{(}E^{\prime}N^{\prime}P_{ft}\bigr{)}$, where $E$ and $E^{\prime}$ are the epochs for pretraining and fine-tuning respectively.
We provide a detailed comparison in Table 5, presenting the total number of parameters, fine-tuned parameters, and the number of epochs used during pretraining and fine-tuning phases. When the ensemble size during pretraining ($N$) is close to the ensemble size used in fine-tuning ($N^{\prime}$), Table 5 shows that the fine-tuning stage requires less than 1% of the pretraining computational time. To quantitatively support this observation, we present detailed computational time comparisons in Table 6. All training times reported here are obtained using a single RTX 4080 Super GPU. The training batch size are adaptively chosen to fully use the 16 GB memory.
Table 5: Comparison of total parameters, fine-tuned parameters, and epochs for various datasets.
@@
Table 6: Comparison of pretraining and fine-tuning computational time for observation noise $\sigma_{y}=1$ with different ensemble sizes on a single RTX 4080 Super GPU.
@@
To further validate the effectiveness of our proposed fine-tuning approach, we will perform additional experiments comparing our method (Subsection 4.4) against alternative fine-tuning strategies. Specifically, we will investigate and quantitatively compare the RMSE performance of the following fine-tuning scenarios: (1) fine-tuning all model parameters and (2) fine-tuning for a longer period (i.e, more epochs).
@@
Figure 6: Comparison of RMSE across training epochs on the Lorenz ’96 model for two fine-tuning strategies under two ensemble sizes, (a) 20 and (b) 40. The vertical dashed red line indicates the stopping epoch of our proposed fine-tuning approach. Results demonstrate that our fine-tuning on a small subset of parameters achieves comparable RMSE performance to fine-tuning all parameters.
@@ 6 Conclusions
In this paper, we propose a novel machine learning-based approach, MNMEF, for state estimation in data assimilation. Our method features a scheme which subsumes the ensemble Kalman filter (EnKF) [8, 9], as a special case, ensuring optimality for linear, Gaussian problems, while extending beyond Gaussian applications through our learnable correction terms. A key advantage of our method is its ability to be trained at one ensemble size and then directly applied, possibly with a cheap fine-tuning of a small parameter subset, for different ensemble sizes. This property is a consequence of the mean field interpretation of the set transformer [44], the machine learning architecture underlying the methodology. Indeed, in this paper we introduce neural operators acting on the metric space of probability measures, which we call measure neural mappings (MNM). We generalize transformers to this setting and show how the set transformer itself may be viewed as a particle approximation of such a mean field model.
@@
Our method adopts a form similar to the Kalman filter in the mean-field perspective, with the core innovation being a parameterized gain matrix analogous to the Kalman gain (Section 2). Since this parameterized gain matrix depends on the current filtering distribution, we employ the set transformer neural network architecture to process such measure inputs (Section 3). In practical applications, our method operates on an ensemble of particles (viewed as an empirical measure) and incorporates learnable mechanisms analogous to inflation and localization techniques in EnKF-based methods, enhancing performance with smaller ensemble sizes (Section 4). Our experiments across multiple challenging systems (Lorenz ’63 [53], Lorenz ’96 [50], and Kuramoto-Sivashinsky [51, 52]) demonstrate that our proposed method consistently outperforms classical approaches including the local ensemble transform Kalman filter (LETKF) [14], with relative improvements of 15-30% for most scenarios (Section 5).
@@
Despite the promising results, our method has limitations and several directions for future improvement. Our current approach is limited by the loss function design, which primarily addresses state estimation rather than general filtering problems. Future work will consider probabilistic loss functions that can be used to estimate the filtering distribution [6, 21].
@@
Additionally, we plan to extend our approach beyond EnKF to include other filtering schemes, such as the ensemble square root filter (ESRF) [47], which is a deterministic version of the EnKF, or variational methods such as iterative EnKF (iEnKF) [15]. The flexibility of our approach stems from the proposed MNM-based learning architecture, which can optimize variables related to the empirical distributions represented by ensembles.
@@
We currently formulate the transformer as a neural operator accepting probability measures as inputs. In future work, we aim to enhance both the architecture and training methodology to simultaneously handle measure and function valued inputs, enabling amortized neural operators that generalize across different dynamics and observation models. Furthermore, since the definition of the attention mechanism is extended to a measure-to-measure operator, we plan to develop a continuous limit analysis examining convergence properties as the number of samples in the empirical distribution approaches infinity. Establishing universal approximation results for this formulation will also be of interest to further strengthen the theoretical foundations of our approach.

### v2

@@ 5.4 Lorenz ’63
Time Step Observation time step: $\Delta t=0.15$; $5$ RK4 integration steps: $\Delta t/5=0.03$
Relative RMSE Refer to caption Refer to caption
Relative Improvement Refer to caption Refer to caption
Lorenz ’63 Ensemble Size Ensemble Size
$\sigma_{y}=0.7$ $\sigma_{y}=1.0$
Figure 7: Comparison results on the Lorenz ’63 system. The upper row of plots shows direct R-RMSE comparisons between different methods, while the lower row illustrates the relative improvement of our fine-tuning method compared to benchmarks. These comparisons are presented for observation noise levels $\sigma_{y}=0.7$ (left column) and $\sigma_{y}=1.0$ (right column). Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain).
@@
In Figure 7, we compare the R-RMSE (5.3) performance of our pretrained MNMEF and fine-tuned MNMEF models against benchmark methods EnKF, ESRF, and IEnKF for the Lorenz ’63 system. The upper row of plots in Figure 7 displays these R-RMSE comparisons for observation noise levels $\sigma_{y}=0.7$ (left plot) and $\sigma_{y}=1.0$ (right plot). The LETKF method is not included in this comparison since the system has only three dimensions, making localization unnecessary. For a fair comparison, we turn off the localization part in our architecture by setting all entries in the localization weight matrices to be $1$. We only fine-tune the parameters $\theta_{\mathrm{gain}}$ for $F^{\mathrm{gain}}$ (4.28) and $\theta_{\mathrm{infl}}$ for $F^{\mathrm{infl}}$ (4.30) for ensemble sizes $N\neq 10$. The lower row of plots in Figure 7 shows the relative improvement $\mathcal{E}_{\mathrm{RI}}$ (5.4) of our fine-tuned MNMEF method as compared to EnKF and IEnKF, for the respective noise levels shown.
@@
The IEnKF [15] uses the same observational information as the standard EnKF, but reformulates the filtering problem as a lag-1 smoothing problem, and solves it iteratively to better handle nonlinearities. While the IEnKF demonstrates excellent performance in low-dimensional systems (e.g., the Lorenz ’63 model, $d=3$), its effectiveness diminishes as dimensionality increases (e.g., the Lorenz ’96 model, $d=40$). This deterioration occurs due to lack of localization in the implementation of the IEnKF we use; however, even with large ensembles with less need for localization, our method outperforms the IEnKF in the higher-dimensional systems. Moreover, IEnKF generally provides a substantial improvement over non-iterative methods only when the observations are sufficiently dense, as shown in [15] using experiments with the Lorenz ’96 model.
@@ 5.5 Robustness To Inherent Randomness
Due to the inherent randomness in both the filtering problem and the algorithm, the same method may exhibit varying performance across different test trajectories. To mitigate this randomness and obtain a more reliable assessment of each method’s effectiveness, we evaluate their performance on $M_{\mathrm{test}}$ trajectories and report the mean R-RMSE $\overline{\mathcal{E}}$ (5.3). While the mean R-RMSE provides a measure of overall accuracy, we are interested in the stability of each method across different trajectories. To quantify this stability, we compute the standard deviation (std) of the R-RMSE values across all test trajectories $\{V_{m}^{\dagger}\}_{m=1}^{M_{\mathrm{test}}}$ and their corresponding estimated trajectories $\{V_{m}\}_{m=1}^{M_{\mathrm{test}}}$, defined as:
$\sigma_{\mathcal{E}}=\sqrt{\frac{1}{M_{\mathrm{test}}}\sum_{m=1}^{M_{\mathrm{test}}}\left(\mathcal{E}(V_{m},V_{m}^{\dagger})-\overline{\mathcal{E}}\right)^{2}}$ (5.8)
@@
Table 4: Standard deviation (std) of the relative root mean square error (R-RMSE) as defined in (5.8). We consider both Pretrain* and Tuned* variants of our MNMEF method. The std shown in the table is an averaged std over ensemble sizes $N=5,10,15,20,40,60,100$. A smaller std indicates better stability of the method. The lowest std in each column is highlighted in bold. The stds for both the pretrained and fine-tuned variants of our MNMEF method are similar and consistently lower than those of the benchmark methods, indicating greater robustness of our approach to the inherent randomness across different test trajectories.
Method Lorenz ’96 (5.5) KS (5.6) Lorenz ’63 (5.7)
$\sigma_{y}=1.0$ $\sigma_{y}=0.7$ $\sigma_{y}=1.0$ $\sigma_{y}=0.7$ $\sigma_{y}=1.0$ $\sigma_{y}=0.7$
Pretrain* 1.03e-2 1.26e-2 8.60e-3 1.07e-2 1.78e-3 1.40e-3
Tuned* 1.08e-2 1.12e-2 8.91e-3 1.07e-2 1.79e-3 1.29e-3
EnKF [8] 2.16e-2 3.10e-2 2.66e-2 3.15e-2 3.41e-3 2.32e-3
ESRF [11] 3.09e-2 1.95e-2 2.59e-2 3.27e-2 2.78e-3 2.01e-3
LETKF [14] 2.63e-2 6.13e-2 2.66e-2 3.58e-2 - -
IEnKF [15] 1.83e-2 2.61e-2 2.23e-2 3.03e-2 2.15e-3 1.44e-3
@@
It is worth noting that in data assimilation literature, metrics similar to variance or standard deviation are often used to measure the spread of ensemble-based methods [6]. However, since our focus in this paper is on state estimation performance, we are not concerned with spread metrics. The stability measure $\sigma_{\mathcal{E}}$ defined in (5.8) is fundamentally different from spread metrics as it quantifies the consistency of estimation errors across different test trajectories rather than the dispersion of ensemble members.
[…]
@@ 5.6 Inflation and Localization
In this section, we will conduct further experiments on learning the inflation (Subsection 4.2.2) and localization (Subsection 4.2.3) components in the architecture of our MNMEF method. These experiments demonstrate that what our method learns has many similarities with what is used in the application of inflation and localization techniques to the benchmark methods.
@@ 5.6.1 Inflation Experiments
In Subsection 4.2.2, we introduce our learning-based inflation scheme (4.31) which depends on the learned output $\widehat{u}^{(n)}_{\theta}$ from the neural network $F^{\mathrm{infl}}$ (4.30).
Relative RMSE Refer to caption Refer to caption
Kuramoto– Ensemble Size Ensemble Size
Sivashinsky $\sigma_{y}=0.7$ $\sigma_{y}=1.0$
Figure 8: Comparison of the trained inflation versus the no inflation scenario when applied to the KS system (5.6), using our MNMEF method, fine-tuned for different ensemble sizes. The plots compare R-RMSE with and without inflation at observation noise levels $\sigma_{y}=0.7$ (left plot) and $\sigma_{y}=1.0$ (right plot). In summary, our trained inflation effectively improves the performance, but even without inflation, our MNMEF does not completely fail.
@@
For our experimental approach, we examine models that are first pretrained and then fine-tuned separately for various ensemble sizes. During the inference process of these models, we manually set the inflation term to zero, i.e. $\widehat{u}^{(n)}_{\theta}=0$, and analyze the resulting changes in the average R-RMSE $\overline{\mathcal{E}}$ (5.3). We conduct these experiments on the KS dynamical system (5.6) only.
Our results are presented in Figure 8. We compare the mean R-RMSE $\overline{\mathcal{E}}$ between our standard architecture with the trained inflation (blue curve) and the no inflation version by setting $\widehat{u}^{(n)}_{\theta}=0$ (red curve) for observation noise $\sigma_{y}=0.7$ and $1.0$. The results in Figure 8 clearly demonstrate the beneficial impact of our trained inflation approach. In addition, our method remains functional, with increased R-RMSE values, when we remove inflation by setting $\widehat{u}^{(n)}_{\theta}=0$ (equivalent to $\alpha=1$ in the EnKF). This addresses our concern, because, if inflation were dominating the architecture by masking the contribution of the state $v^{(n)}$, removing it would cause model collapse.
@@ 5.6.2 Localization Experiments
In this subsection, we present experimental comparisons between our learned function $g_{\theta}$ and the Gaspari–Cohn (GC) function with the optimal localization radius, found by grid-search, for LETKF; see Figure 9 . For ensemble sizes $N=5,10,15,20,40,60,100$, the optimal GC radius values chosen are $1,1,2,3,4,4,6$ (multiplied by $\sqrt{10/3}$ when implemented [48]), respectively. The learned function $g_{\theta}$ (shown as red curves with the mean and standard deviation of the learned $g_{\theta}$ across assimilation time steps) demonstrates a similar shape to the GC function with optimally chosen radius (blue curves). Notably, our approach dynamically adjusts the localization weights at each assimilation step based on ensemble size and distribution, while the GC function remains static throughout the assimilation process.
$\sigma_{y}=0.7$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\sigma_{y}=1.0$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$N$ 5 10 15 20 40 60 100
@@
The results in Figure 9 demostrates the effectiveness of our learned localization. Despite imposing no explicit constraints on shape or monotonicity, the learned function $g_{\theta}$ naturally exhibits similar behavior to the GC function: it decreases monotonically with distance and shares a similar shape. A key difference is that our function occasionally exceeds values of 1 at short distances, reflecting a combined localization and inflation effect where covariance between closely spaced dimensions is amplified, since we are learning the inflation and localization simultaneously. This adaptivity allows $g_{\theta}$ to dynamically adjust to the ensemble distribution, leading to improved performance over the fixed GC function, as supported by RMSE results in Subsections 5.2 and 5.3.
@@ 5.7 Fine-Tuning for Different Ensemble Size
Firstly, we consider the time complexity between the pretraining stage and fine-tuning stage. Let $P_{\mathrm{pretrain}}$ and $P_{\mathrm{ft}}$ denote the total number of pretraining and fine-tuning parameters respectively, i.e.
$P_{\mathrm{pretrain}}$ $=|\theta_{\mathrm{ST}}|+|\theta_{\mathrm{gain}}|+|\theta_{\mathrm{infl}}|+|\theta_{\mathrm{loc}}|,$ (5.9a)
$P_{\mathrm{ft}}$ $=|\theta_{\mathrm{gain}}|+|\theta_{\mathrm{infl}}|+|\theta_{\mathrm{loc}}|.$ (5.9b)
For fixed dynamic and training trajectories, the time complexity for pretraining with the ensemble size $N$ is $\mathcal{O}\bigl{(}ENP_{t}\bigr{)}$ while the fine-tuning with the ensemble size $N^{\prime}$ has the time complexity $\mathcal{O}\bigl{(}E^{\prime}N^{\prime}P_{ft}\bigr{)}$, where $E$ and $E^{\prime}$ are the epochs for pretraining and fine-tuning, respectively.
We provide a detailed comparison in Table 5, presenting the total number of pretraining parameters $P_{\mathrm{pretrain}}$, fine-tuning parameters $P_{\mathrm{ft}}$, and the number of epochs $E$ used during the pretraining and fine-tuning phases. When the ensemble size during pretraining ($N$) is close to the ensemble size used in fine-tuning ($N^{\prime}$), Table 5 shows that the fine-tuning stage requires less than 1% of the pretraining computational time. To quantitatively support this observation, we present detailed computational time comparisons in Table 6. All training times reported here are obtained using a single RTX 4080 Super GPU. The training batch sizes are adaptively chosen to fully use the 16 GB memory.
Table 5: Comparison of total parameters, fine-tuning parameters, and epochs for various datasets.
@@
Table 6: Comparison of pretraining and fine-tuning computational time for observation noise $\sigma_{y}=1.0$ with different ensemble sizes on a single RTX 4080 Super GPU. The ratio is between the computation time of the fine-tuning stage and the time of the pretraining stage. The fine-tuning time increases with the ensemble size, but it remains significantly faster compared to the pretraining stage.
@@
Figure 10: Comparison of R-RMSE across training epochs on the Lorenz ’96 model for two fine-tuning strategies under two ensemble sizes, (a) 20 and (b) 40. The vertical dashed red line indicates the stopping epoch of our proposed fine-tuning approach. Results demonstrate that our fine-tuning on a small subset of parameters achieves comparable RMSE performance to fine-tuning all parameters.
To further validate the effectiveness of our proposed fine-tuning approach, we perform additional experiments comparing our method (Subsection 4.4) against alternative fine-tuning strategies. Specifically, we investigate and quantitatively compare the RMSE performance of the following fine-tuning scenarios: (1) fine-tuning all model parameters and (2) fine-tuning for a longer period (i.e., more epochs).
@@ 6 Conclusions
In this paper, we propose a novel machine learning-based approach, MNMEF, for state estimation in data assimilation. Our method features a scheme which subsumes the ensemble Kalman filter (EnKF), as a special case, ensuring optimality for linear, Gaussian problems, while extending beyond Gaussian applications through our learnable correction terms. A key advantage of our method is its ability to be trained at one ensemble size and then directly applied, possibly with a cheap fine-tuning of a small subset of parameters, for different ensemble sizes. This property is a consequence of the mean-field interpretation of the set transformer, the machine learning architecture underlying the methodology. Indeed, in this paper we introduce neural operators acting on the metric space of probability measures, which we call measure neural mappings (MNM). We generalize transformers to this setting and show how the set transformer itself may be viewed as a particle approximation of such a mean-field model.
@@
Our method adopts a form similar to the Kalman filter in the mean-field perspective, with the core innovation being a parameterized gain matrix analogous to the Kalman gain (Section 2). Since this parameterized gain matrix depends on the current filtering distribution, we employ the set transformer neural network architecture to process such measure inputs (Section 3). In practical applications, our method operates on an ensemble of particles (viewed as an empirical measure) and incorporates learnable mechanisms analogous to inflation and localization techniques in EnKF-based methods, enhancing performance with smaller ensemble sizes (Section 4). Our experiments across multiple challenging systems (Lorenz ’63, Lorenz ’96, and Kuramoto-Sivashinsky) demonstrate that our proposed method consistently outperforms classical approaches including the local ensemble transform Kalman filter (LETKF), with relative improvements of 15-30% for most scenarios (Section 5).
@@
Despite the promising results, our method has limitations and several directions for future improvement. Our current approach is limited by the loss function design, which primarily addresses state estimation rather than general filtering problems. Future work will consider probabilistic loss functions that can be used to estimate the filtering distribution.
@@
Additionally, we plan to extend our approach beyond EnKF to include other filtering schemes, such as the ensemble square root filter (ESRF), which is a deterministic version of the EnKF, or variational methods such as iterative EnKF (iEnKF). The flexibility of our approach stems from the proposed MNM-based learning architecture, which can optimize variables related to the empirical distributions represented by ensembles.
@@
We currently formulate the transformer as a neural operator accepting probability measures as inputs. In future work, we aim to enhance both the architecture and training methodology to simultaneously handle measure and function valued inputs, enabling amortized neural operators that generalize across different dynamics and observation models. Establishing universal approximation results and statistical guarantees for this formulation will also be of interest to further strengthen the theoretical foundations of our approach.